OpenStudy (ovidia):

-20x-{3x+[16x-(-9x+8)]} @jango_IN_DTOWN

OpenStudy (ovidia):

highest to lowest power of the variable

OpenStudy (jango_in_dtown):

-20x-{3x+[16x-(-9x+8)]}= -20x-[3x+16x+9x-8]=20x-(28x-8)=8-8x i.e. -8x+8

OpenStudy (ovidia):

so ? -8x+8? @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

CORRECT

OpenStudy (ovidia):

it says it was wrong thanx tho @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

-20x-{3x+[16x-(-9x+8)]}= -20x-[3x+16x+9x-8]=-20x-(28x-8)==-48X+8